# Investigating Mass Segregation of the Binary Stars in the Open Cluster NGC 6819

### v2

@@ 1 Introduction
NGC 6819 is an intermediate-aged open cluster and home to a rich binary population (Hole et al., 2009; Milliman et al., 2014; Chen et al., 2022). With an age of about 2.5 Gyr (e.g., Kalirai et al., 2001; Basu et al., 2011; Jeffries et al., 2013; Sandquist et al., 2013; Yang et al., 2013a; Ak et al., 2016; Brewer et al., 2016), NGC 6819 has persisted through 5-10 half-mass relaxation times (e.g., Kalirai et al., 2001; Kang & Ann, 2002; Karataş et al., 2023) and therefore should be relatively dynamically relaxed. One product of two-body relaxation is an expectation that the more massive single and binary stars in the cluster will occupy a more centrally concentrated spatial distribution than those with lower mass (e.g., King, 1962a; Giersz & Heggie, 1997).
Several previous studies have found evidence that NGC 6819 is mass segregated. Kalirai et al. (2001) compared luminosity functions in radial annuli and found the central annulus to be weighted toward brighter (higher-mass) stars as compared to an almost flat luminosity function observed for the entire cluster. They interpret this finding as evidence for dynamical evolution causing higher mass stars to sink to the inner regions of the cluster. Kalirai et al. (2001) also examined mass functions across eight different annuli, finding that the slope changes from positive (more higher-mass stars) to negative (more lower-mass stars) with increasing radial distance from the cluster center, consistent with the expectations of dynamical evolution. Kang & Ann (2002) constructed cumulative distribution functions (CDFs) of the number of stars with respect to the distance from the cluster center in bins of magnitude, finding that brighter stars have more centrally concentrated radial distributions. They also calculated “half number radii” for their brightest and faintest magnitude bins confirming this result. Yang et al. (2013b) also constructed mass functions in radial annuli and confirm the cluster to be mass segregated.
@@ 2 Cluster Membership, Global Parameters and our Primary Stellar Sample
We follow a nearly identical procedure to Childs et al. (2024) to prepare a sample for analysis of mass segregation in NGC 6819. In short, first we download Gaia (Gaia Collaboration et al., 2023), Pan-STARRS (Magnier et al., 2020) and 2MASS (Skrutskie et al., 2006) data for all stars within the cluster’s effective radius of (Childs et al. 2024, equivalent to 4.5 core radii as derived from a King 1962a model fit to our data) from the center of NGC 6819, = and = (Platais et al., 2013). We use the Gaia DR3 radial-velocity, proper-motion and parallax distributions to derive priors on the cluster membership for each star (a given star’s membership prior is calculated based on the distance that the star’s Gaia kinematics and parallax values are measured away from the cluster’s mean measurements; see Childs et al. 2024 for more details). We follow the same procedure as Childs et al. (2024) to use the Bayesian Analysis for Stellar Evolution with Nine Parameters (BASE-9) statistics software suite (von Hippel et al., 2006; van Dyk et al., 2009; Robinson et al., 2016) and the PARSEC isochrone models (Bressan et al., 2012) to derive the posterior distributions for the global cluster parameters (cluster age, distance, reddening and metallicity) for NGC 6819. Childs et al. 2024 find the following median values and 1 uncertainties for the posterior distributions of each global cluster parameter: an age of Gyr, a distance of kpc, [Fe/H] = dex, and mmag. We show isochrones using these median values in Figure 1.
@@ 3.1 Relaxation timescales for NGC 6819
where is the number of cluster members, is a constant we assume to be 0.1 (Giersz & Heggie, 1994), is the half-mass radius, is the total mass of the cluster, and is the gravitational constant. To estimate the half-mass relaxation time, we use all cluster members in our sample, no longer limited by the primary sample selected above. We estimate a half-mass radius both by investigating the radial mass distribution (which we assume is incomplete) and using a simple conversion from core radius (Heggie & Hut, 2003), and find to be between roughly 5 and 10 pc, assuming a distance of kpc and accounting for projection effects (by multiplying our observed projected half-mass radius by a factor of 4/3, following Spitzer 1987). For our calculation we use pc (de-projected). Within our sample we find and out to the cluster’s effective radius. Previous literature reports NGC 6819 having a mass and number of members in the range and (Kalirai et al., 2001; Kang & Ann, 2002; Yang et al., 2013b). Due to the varying methodologies employed to calculate the total cluster mass, each incomplete in its own way, we opt for a cautious approach, settling on an intermediate value of M = 2700 600 and N = 2400 400. This results in = 440 230 Myr. This is somewhat larger than, though still consistent with previous values from the literature (Kalirai et al., 2001; Kang & Ann, 2002; Yang et al., 2013b). Given the age from our BASE-9 analysis of Gyr, we find that NGC 6819 has survived more than five half-mass relaxation times.
@@ 5 Discussion and Conclusions
Through our analysis above, using a sample of only confident kinematic, spatial and photometric cluster members, we verify that NGC 6819 is mass segregated. These results confirm findings in the literature (e.g., Kalirai et al., 2001; Kang & Ann, 2002; Yang et al., 2013b; Karataş et al., 2023) using somewhat different samples and different techniques. Importantly, here for the first time we also find that the binary stars show strong evidence of mass segregation, both with respect to the single stars and within the binary population itself.

### v3

@@ 1 Introduction
NGC 6819 is an intermediate-aged open cluster and home to a rich binary population (Hole et al., 2009; Milliman et al., 2014; Chen et al., 2022). With an age of about 2.5 Gyr (e.g., Kalirai et al., 2001; Basu et al., 2011; Jeffries et al., 2013; Sandquist et al., 2013; Yang et al., 2013; Ak et al., 2016; Brewer et al., 2016), NGC 6819 has persisted through 5-10 half-mass relaxation times (e.g., Kalirai et al., 2001; Kang & Ann, 2002; Karataş et al., 2023) and therefore should be relatively dynamically relaxed. One product of two-body relaxation is an expectation that the more massive single and binary stars in the cluster will occupy a more centrally concentrated spatial distribution than those with lower mass (e.g., King, 1962; Giersz & Heggie, 1997).
Several previous studies have found evidence that NGC 6819 is mass segregated. Kalirai et al. (2001) compared luminosity functions in radial annuli and found the central annulus to be weighted toward brighter (higher-mass) stars as compared to an almost flat luminosity function observed for the entire cluster. They interpret this finding as evidence for dynamical evolution causing higher mass stars to sink to the inner regions of the cluster. Kalirai et al. (2001) also examined mass functions across eight different annuli, finding that the slope changes from positive (more higher-mass stars) to negative (more lower-mass stars) with increasing radial distance from the cluster center, consistent with the expectations of dynamical evolution. Kang & Ann (2002) constructed cumulative distribution functions (CDFs) of the number of stars with respect to the distance from the cluster center in bins of magnitude, finding that brighter stars have more centrally concentrated radial distributions. They also calculated “half number radii” for their brightest and faintest magnitude bins confirming this result. Yang et al. (2013) also constructed mass functions in radial annuli and confirm the cluster to be mass segregated.
@@ 2 Cluster Membership, Global Parameters and our Primary Stellar Sample
We follow a nearly identical procedure to Childs et al. (2024) to prepare a sample for analysis of mass segregation in NGC 6819. In short, first we download Gaia (Gaia Collaboration et al., 2023), Pan-STARRS (Magnier et al., 2020) and 2MASS (Skrutskie et al., 2006) data for all stars within the cluster’s effective radius of (Childs et al. 2024, equivalent to 4.5 core radii as derived from a 1962AJ.....67..471K model fit to our data) from the center of NGC 6819, = and = (Platais et al., 2013). We use the Gaia DR3 radial-velocity, proper-motion and parallax distributions to derive priors on the cluster membership for each star (a given star’s membership prior is calculated based on the distance that the star’s Gaia kinematics and parallax values are measured away from the cluster’s mean measurements; see Childs et al. 2024 for more details). We follow the same procedure as Childs et al. (2024) to use the Bayesian Analysis for Stellar Evolution with Nine Parameters (BASE-9) statistics software suite (von Hippel et al., 2006; van Dyk et al., 2009; Robinson et al., 2016) and the PARSEC isochrone models (Bressan et al., 2012) to derive the posterior distributions for the global cluster parameters (cluster age, distance, reddening and metallicity) for NGC 6819. Childs et al. 2024 find the following median values and 1 uncertainties for the posterior distributions of each global cluster parameter: an age of Gyr, a distance of kpc, [Fe/H] = dex, and mmag. We show isochrones using these median values in Figure 1.
@@ 3.1 Relaxation timescales for NGC 6819
where is the number of cluster members, is a constant we assume to be 0.1 (Giersz & Heggie, 1994), is the half-mass radius, is the total mass of the cluster, and is the gravitational constant. To estimate the half-mass relaxation time, we use all cluster members in our sample, no longer limited by the primary sample selected above. We estimate a half-mass radius both by investigating the radial mass distribution (which we assume is incomplete) and using a simple conversion from core radius (Heggie & Hut, 2003), and find to be between roughly 5 and 10 pc, assuming a distance of kpc and accounting for projection effects (by multiplying our observed projected half-mass radius by a factor of 4/3, following Spitzer 1987). For our calculation we use pc (de-projected). Within our sample we find and out to the cluster’s effective radius. Previous literature reports NGC 6819 having a mass and number of members in the range and (Kalirai et al., 2001; Kang & Ann, 2002; Yang et al., 2013). Due to the varying methodologies employed to calculate the total cluster mass, each incomplete in its own way, we opt for a cautious approach, settling on an intermediate value of M = 2700 600 and N = 2400 400. This results in = 440 230 Myr. This is somewhat larger than, though still consistent with previous values from the literature (Kalirai et al., 2001; Kang & Ann, 2002; Yang et al., 2013). Given the age from our BASE-9 analysis of Gyr, we find that NGC 6819 has survived more than five half-mass relaxation times.
@@ 5 Discussion and Conclusions
Through our analysis above, using a sample of only confident kinematic, spatial and photometric cluster members, we verify that NGC 6819 is mass segregated. These results confirm findings in the literature (e.g., Kalirai et al., 2001; Kang & Ann, 2002; Yang et al., 2013; Karataş et al., 2023) using somewhat different samples and different techniques. Importantly, here for the first time we also find that the binary stars show strong evidence of mass segregation, both with respect to the single stars and within the binary population itself.